OpenStudy (anonymous):

You are given functions f and g such that f(n)=O(g(n)). Is f(n)∗log2(f(n)c)=O(g(n)∗log2(g(n))) ? (Here c is some constant >0. You can assume that f and g are always bigger than 1. a) true b)false c)depends on c d) depends on f and g

OpenStudy (anonymous):

I think could go this way: f(n)=O(g(n)) means |f(n)|<= M |g(n)| or |f(n)|/|g(n)| <= M so you have to prove, given that f(n)=O(g(n)), that |f(n)∗log2(f(n)c)|/|g(n)∗log2(g(n))| <= N |f(n)∗log2(f(n)c)|/|g(n)∗log2(g(n))| <= M|log2f(n)c)|/|log2(g(n))| = =M log2|(f(n)c -g(n))| <= M log2 (|M|g(n)|c - g(n)|). So i think it depends on c.

OpenStudy (anonymous):

thanks myko was really stuck on this

OpenStudy (anonymous):

not 100% sure , :). Just an idea

OpenStudy (anonymous):

well sumthing is better than none

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

for two (positive) functions f and g such that f(n)=O(g(n)). Is 2f(n)=O(2g(n)) ? (Multiple answers may be correct.) Sometimes Never Yes if f(n)≤g(n) for all sufficiently large n Always

OpenStudy (anonymous):

if i am not wrong for it the answer should be Yes if f(n)≤g(n) for all sufficiently large n

OpenStudy (anonymous):

2|f(n)|/2|g(n)| =|f(n)|/|g(n)|<= M so always, i would say. By the way you can take out the ||, since functions always positive

OpenStudy (anonymous):

Is 2f(n)=O(2g(n)) ? (

OpenStudy (anonymous):

Always

OpenStudy (anonymous):

actually it is 0(2powg(n))

OpenStudy (anonymous):

if f(n)=O(g(n)), then 2f(n)=O(2g(n))

OpenStudy (anonymous):

ai get that but the qquestion is 2f(n)=O(2powerg(n))

OpenStudy (anonymous):

you mean? \[2f(n) =O(2^{g(n)})\]